OpenStudy (anonymous):

A recipe asks for 2 1/3 cup of flour. How many cups of flour are needed for 3 1/2 recipes? A.5 5/6 cups B.8 1/3 cups C.6 1/6 cups D.8 1/6 cups

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

C ?

OpenStudy (anonymous):

2 1/3 Is needed so 2+2+2= 6 1/6 idk lol

OpenStudy (anonymous):

@AloneS is way better at this than me

OpenStudy (anonymous):

2 1/3 x 3 1/2 First we change the fractions from mixed fractions to top heavy ones. 2 1/3 = 7/3 3 1/2 = 7/2 Then, we multiply together 7/3 x 7/2 = 49/6 Finally convert back into a mixed fraction 8 1/6 D.

OpenStudy (anonymous):

omg lol